OpenStudy (kaylaprincess):

Can I get help on how to start solving this, along with the formula? Studying for Midterm tomorrow. Aimee currently has an account balance of $2,199.21. She opened the account 9 years ago with a deposit of $1,668.17. If the interest compounds quarterly, what is the interest rate on the account?

OpenStudy (kaylaprincess):

@zepdrix

OpenStudy (kaylaprincess):

@Awolflover1 @samanthagreer @quickstudent

OpenStudy (samanthagreer):

What do you think it is

OpenStudy (kaylaprincess):

The formula?

OpenStudy (kaylaprincess):

or?

OpenStudy (samanthagreer):

the anwer

OpenStudy (kaylaprincess):

i don't know...

OpenStudy (kaylaprincess):

I don't even know what formula to use.

OpenStudy (kaylaprincess):

Anything appreciated.

zepdrix (zepdrix):

\[\large\rm A=P\left(1+\frac{r}{n}\right)^{nt}\] P is our principle or starting amount r is the interest rate n is the number of compoundings per period t is the period (number of years in this case) A is the amount we end up with.

zepdrix (zepdrix):

They gave us a bunch of the pieces, and they expect us to solve for the unknown piece, which is r in this case.

OpenStudy (kaylaprincess):

in this situation, just for personal preference, can we use x?

zepdrix (zepdrix):

For your unknown? Sure!\[\large\rm A=P\left(1+\frac{x}{n}\right)^{nt}\]

OpenStudy (kaylaprincess):

okay lemme plug everything in and try to solve, then can i check back?

zepdrix (zepdrix):

Do you understand the value of each part that they gave us? t = 9 years P = 1668.17 n is a little tricky. quarterly means we'll be calculating it 4 times a year, ya?

OpenStudy (kaylaprincess):

Oh!!!! I tried to solve this with that formula!

OpenStudy (kaylaprincess):

yes i do!

OpenStudy (kaylaprincess):

correct!

OpenStudy (kaylaprincess):

the problem i came to was when we get \[2,199.21=1668.17(1+\frac{ x }{ 4 })^{36}\]

OpenStudy (kaylaprincess):

the \[\frac{ x }{ 4 }\] How do i deal with that thing?

OpenStudy (kaylaprincess):

@zepdrix

zepdrix (zepdrix):

Ahh sorry was making a sammich >.<

OpenStudy (kaylaprincess):

no youre fine! Just checkn!

zepdrix (zepdrix):

\[\large\rm 2,199.21=1668.17\left(1+\frac{x}{4}\right)^{36}\]Hmm bunch of weird stuff being applied to our x. We'll have to undo all of this, one step at a time. Notice this big block of stuff is being `multiplied` by 1668.17. So we'll have to `divide` both sides by this value to undo that step.

OpenStudy (kaylaprincess):

Because of order and operations duh!

OpenStudy (kaylaprincess):

i should've know that to begin lol

zepdrix (zepdrix):

\[\large\rm \frac{2199.21}{1668.17}=\left(1+\frac{x}{4}\right)^{36}\]Ok with that step?

OpenStudy (kaylaprincess):

1.3183?

zepdrix (zepdrix):

\[\large\rm 1.318=\left(1+\frac{x}{4}\right)^{36}\]Ok. So the `division` took care of that `multiplication`. Now we need to somehow deal with this `exponent`. Do you remember how to get rid of an exponent?

OpenStudy (kaylaprincess):

oh god, i dont :-(

zepdrix (zepdrix):

We undo `powers` using `roots`. For example if we had an equation like this: \(\large\rm x^2=4\) We would undo the `square` on the x by taking a `square root` of each side.\[\large\rm \sqrt{x^2}=\sqrt4\]\[\large\rm x=\pm2\]

zepdrix (zepdrix):

We have a 36th power, so we'll need to take a 36th root of each side.

OpenStudy (kaylaprincess):

im horrible at roots, but i think i remember some things about htat stuff. I wouldn't know how to continue from here though

zepdrix (zepdrix):

\[\large\rm \sqrt[36]{1.318}=\left(1+\frac{x}{4}\right)\]So we'll take a 36th root of each side, that will undo the 36 power on the right side of the equation.

zepdrix (zepdrix):

If you use your calculator, you should get something like 1.007699 on the left side of the equation. Maybe we'll round that to 1.008

zepdrix (zepdrix):

\[\large\rm 1.008=\left(1+\frac{x}{4}\right)\]

OpenStudy (kaylaprincess):

could we solve without roots?

OpenStudy (kaylaprincess):

i think i have the answer though : 3.1% ?

zepdrix (zepdrix):

I'm coming up with x=0.032 which is 3.2% (But that was with the rounded values, so maybe yours is more accurate).

zepdrix (zepdrix):

without roots? Hmm I don't think so :o

zepdrix (zepdrix):

Yayyy team, we did it! :)

OpenStudy (kaylaprincess):

Thank you for your help!